OpenStudy (nick88888888):

(a + b) - c = a + (b - c) errm use properties and definitions of operations to show that the statement is true. I just want to know how to do the steps :c Or Justify each step

OpenStudy (amistre64):

we could just add c to both sides i spose

OpenStudy (amistre64):

distrbute the 1 thru the paranthesis, and add a c to both sides ... is my thoughts

OpenStudy (nick88888888):

@amistre64 errr theres no 1....theres only a, b, and c...no numbers .-. and that's not how you do it lul

OpenStudy (amistre64):

there is a 1, its implied

OpenStudy (amistre64):

if you have another way to approach it, then by all means, go for it

OpenStudy (nick88888888):

Heres a example "with number" 7a + (4+5a) = 12a + 4 7a + (4+5a) = 7a + (5a+4) = commutative property of addition = (7a + 5a) + 4 = associative property of addition = 12a + 4 = combine like terms @amistre64

OpenStudy (amistre64):

are distribution and inverse part of your properties?

OpenStudy (nick88888888):

@amistre64 Yes.

OpenStudy (amistre64):

hmm, they and it seems that you are not really equating both sides, but rather trying to transform the left into the right ....

OpenStudy (amistre64):

well, (a+b) - c 1(a+b)-c 1a + 1b - c a + b - c a + b + (- c) a + 1b + 1(- c) a + 1(b + (- c)) there are a myriad of ways to approach this. but the simplest just seems to be associativity of addition

OpenStudy (amistre64):

associativity is an application of the distributive property ... and combining like terms is undoing the distributive property: 7a + 4a = (7+4)a, by factoring out an a